OpenStudy (anonymous):

pls help solve these!! 1. y=7/(2x+48) 2. y=1/(x-2) -3

OpenStudy (anonymous):

Are you supposed to solve for x?

OpenStudy (anonymous):

what are we suppose to solve for?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

y=7/(2x/+48) , why is there a dividion sign iside the parenthese?

OpenStudy (anonymous):

there isnt one... typo my bad lol i fixed it

jimthompson5910 (jim_thompson5910):

y=7/(2x+48) y(2x+48) = 7 2xy + 48y = 7 2xy = 7 - 48y x = (7 - 48y)/(2y) So the solution is \[\Large x = \frac{7-48y}{2y}\]

OpenStudy (anonymous):

what are the excluded values?

jimthompson5910 (jim_thompson5910):

y cant be zero (to avoid division by zero)